True or false p lies in plane b the line containing p and q must lie in plane b?
False.
In order for the line PQ to lie in plane B, then both P and Q must lie in plane B.
How many face cards are there in a traditional deck of playing cards?
J, Q, K of four suits = 12 face cards.
If two numbers are complementary what is their probability?
On the basis that numbers are a continuous variable, the probability of any particular number (or pair) is 0.
How many outcomes can you get if you make seven number random phone numbers?
The answer depends on what you are measuring. The random variable could be:
Other variables will generate other outcomes.
What is the probability of a 5 not being rolled on a dice?
Probability is desired options over total options.
There are 6 faces on a standard dice, so NOT rolling a 5 is 5/6.
When three dice are thrown together what is the probability of getting a total score of 10?
As there are 3 dice, it is quite easy (and quick) to work out the possible ways a sum of 10 can be made:
If the first die is a 1, the second and third dice could be {3, 6}; {4, 5}; {5, 4}; {6, 3} - 4 possibilities
If the first die is a 2, the second and third dice could be {2, 6}; {3, 5}; ...; {6, 2} - 5 possibilities
If the first die is a 3, the second and third dice could be {1, 6}; {2, 5}; ...; {6, 1} - 6 possibilities
If the first die is a 4, the second and third dice could be {1, 5}; {2, 4}; ...; {5, 1} - 5 possibilities
If the first die is a 5, the second and third dice could be {1, 4}; {2, 3}; ...; {4, 1} - 4 possibilities
If the first die is a 6, the second and third dice could be {1, 3}; {2, 2}; {3, 1} - 3 possibilities
So there are 4 + 5 + 6 + 5 + 4 + 3 = 27 possible ways of getting a sum of 10
With 3 dice there are 6 × 6 × 6 = 216 possible ways of them landing
→ pr(sum 10) = 27/216 = 1/8 = 0.125 = 12.5%
How is probability used in genetics?
probability is used in genetics to determine the possibilities of offspring having a certain trait
2 of hearts, 2 of diamonds, 3 of diamonds, and 4 of hearts
4/8 x 3/5
When two dice are rolled find the probability of getting a sum less than 13?
Assuming these are regular dice, the probability is 1.
What is the probabililty of at least 2 people same birthday from a group of 13 people?
19.4%
CALCULATION:
The probability of at least 2 people having the same birthday in a group of 13
people is equal to one minus the probability of non of the 13 people having the
same birthday.
Now, lets estimate the probability of non of the 13 people having the same birthday.
(We will not consider 'leap year' for simplicity, plus it's effect on result is minimum)
1. We select the 1st person. Good!.
2. We select the 2nd person. The probability that he doesn't share the same
birthday with the 1st person is: 364/365.
3. We select the 3rd person. The probability that he doesn't share the same
birthday with 1st and 2nd persons given that the 1st and 2nd don't share the same
birthday is: 363/365.
4. And so forth until we select the 13th person. The probability that he doesn't
share birthday with the previous 12 persons given that they also don't share
birthdays among them is: 353/365.
5. Then the probability that non of the 13 people share birthdays is:
P(non of 13 share bd) = (364/365)(363/365)(362/365)∙∙∙(354/365)(353/365)
P(non of 13 share bd) ≈ 0.805589724...
Finally, the probability that at least 2 people share a birthday in a group of 13
people is ≈ 1 - 0.80558... ≈ 0.194 ≈ 19.4%
The above expression can be generalized to give the probability of at least x =2
people sharing a birthday in a group of n people as:
P(x≥2,n) = 1 - (1/365)n [365!/(365-n)!]
80%
it is a full cover bet is any bet which consists of all available multiple bets over a given number of selections.
What is meant by the term probability?
Probability means - The likelihood that a particular event will occur.
What is the probability of picking a queen from a pack of playing cards?
What is the probability of choosing something to wear?
depends on choices. 4 pair of shoes -- pick 1. 6 shirts pick 1. etc
1/4 x 1/6 x etc
What is the difference between 'probability' and 'proportion'?
One in six. One in two.